OpenStudy (anonymous):

if annual growth rate is 6.5% what is the equivalent continuous growth rate?

OpenStudy (anonymous):

with annual rate 0.065, after one year (with only one compound), you will get P(1 + 0.065). Equal this with the amount that you will get when you compound continuously, that is, P(e^r). This gives 1.065 = e^r ln(1.065) = r